# Drop rebound at low Weber number

### v1

@@ 2.2 Experimental protocol
To begin a set of experiments, a reservoir connected to the drop generator is filled with silicone oil, and the system is manually purged to remove any trapped air (Harris et al., 2015). Next, the pulse width controlling the piezoelectric disks actuation is tuned to produce a single drop with minimal surface oscillations at impact. These settings vary based on the working fluid and nozzle. Next, the nozzle is lowered until it rests just above the substrate, and the system height is set to zero. The nozzle is raised to the desired initial drop height, typically 1 or 2 mm, and a droplet is generated. The drop falls under the action of gravity and its motion is recorded from release until bouncing ceases. After coming to rest, an external pulse of air is applied and sweeps the drop from the substrate before it can coalesce with the film, ensuring the surface remains pristine for subsequent experiments. The nozzle height is raised incrementally, and three to eight drops are recorded at each height. The experiments conclude once impact stresses become strong enough to pierce the intervening air layer (Hao et al., 2015), which occurs at $W\kern-1.49994pte\,\approx 10$ for all $O\kern-0.50003pth\,$, or drop retraction leads to droplet ejection for low $O\kern-0.50003pth\,$ drops (Richard & Quéré, 2000; Bartolo et al., 2006), as shown in supplemental movie 2.
@@ 3.1.1 Drop motion
Here, $P_{l}$ are Legendre polynomials, and $\mathscr{A^{\prime}}_{l}$ denotes the time-dependent amplitude of the $l^{\text{th}}$ mode. The $l=0$ mode is excluded due to incompressibility. The $l=1$ mode is also excluded in the present formulation due to the choice of the drop’s center of mass as the origin, since the $l=1$ mode shifts the center of mass. Figure 3($c$) demonstrates the synthesis of Legendre polynomials for modes $l=2-4$ into a sample deformed drop shape. Similarly, the contact pressure $p^{\prime}$ exerted by the substrate on the drop is expressed as
@@
Figure 3: ($a$) Schematic of an axisymmetric drop impacting a rigid substrate, where $\zeta^{\prime}\left(\theta,t\right)$ is the dimensional interface, $h^{\prime}\left(t\right)$ the height of the center of mass of the drop, and $S\left(t\right)$ the contact surface. ($b$) Angular discretization of the right hemisphere of an axisymmetric drop during impact. Filled circles represent mesh points where the drop and substrate touch, while hollow circles denote points along the liquid-gas interface. ($c$) An example of shape reconstruction through Legendre synthesis for modes $l=2-4$.
@@ 3.1.3 Non-dimensionalization
[…]
@@ 3.2 Direct numerical simulation
We develop a computational framework based on the Basilisk open-source code structure (Popinet, 2009, 2015) as a comparative benchmark for the mathematical model described in Section 3.1, and to interrogate quantities difficult to visualize experimentally. The fully nonlinear implementation is used to resolve the time-dependent Navier-Stokes equations in both liquid and gas phases, with physical properties of the latter taken to be the values of air at room temperature i.e. $\rho_{g}=0.00121$ g cm-3 and $\mu_{g}=0.000181$ g cm-1 s-1, in an attempt to replicate the experimental conditions. The adaptive mesh refinement and parallelization capabilities of the implementation make it a particularly good candidate for this multi-scale impact problem. The code base for bouncing problems has been constructed and systematically validated for several years in our extended group, covering fluid-structure interaction (Galeano-Rios et al., 2021) and liquid-liquid impact (Alventosa et al., 2023), with comprehensive numerical campaigns in the wider community (Sanjay et al., 2023a, b) providing key insights into related regimes.
@@
Figure 4: ($a$) Visualization of the computational domain comprising of liquid (blue) and gas regions (white), and underlying adaptive grid cell structure in the direct numerical simulation, for a typical $W\kern-1.34995pte\,=0.253$ impact taken at the point of maximal radius deformation $t=0.746$ ms. ($b$) Dimensionless distance above the substrate $\bar{z}=z/R$ against time $\bar{t}=t/t_{\sigma}$ for the top, bottom, and center of mass of a drop for $W\kern-1.34995pte\,=0.253$, $B\kern-0.45003pto\,=0.02$ and $O\kern-0.45003pth\,=0.03$. Black solid lines represents experiment, while blue solid lines and yellow dashed lines represent the KM model and direct numerical simulations, respectively. ($c$) Overlay of the predicted shape from the KM model (solid blue) and direct numerical simulations (dashed yellow) with experimental images corresponding to ($b$). The drop radius is $R=0.203$ mm and the time interval between images is 0.375 ms.
The axisymmetric computational domain set up in cylindrical coordinates measures $8R$ across both ($r-$ and $z-$) spatial dimensions. To ensure sufficient resources are dedicated to the delicate bouncing dynamics, we use a strong grid cell refinement level with a minimal grid cell size of approximately $1/20\ \mu$m, verified to be robust across our parameter regime of interest. This stringent resolution specification was particularly relevant at the lowest Weber numbers studied herein, where the implementation requires most resilience in view of increased numerical stiffness. Adaptivity was tailored around the evolving location of the drop interface and changes in magnitude of the velocity components in the flow, as illustrated in figure 4($a$). Special care was taken during runtime via stringent projection solver tolerances and increased number of internal iterations set for the multigrid cycle to ensure a high degree of accuracy. Furthermore, a zero Dirichlet boundary condition for the color function was prescribed at the level of the solid surface to prevent touchdown and contact, with all tests resulting in the drop rebounding or resting on the surface. We employ a criterion based on the drop interface crossing a threshold level of $0.02R$ above the substrate to define the initialization and end of contact as discussed in Appendix A. This criterion enables us to build a consistent comparative framework with the model and experiment. Under these specifications, a typical run is characterised by $\mathcal{O}(10^{5})$ grid cells and $\mathcal{O}(10-100)$ CPU hours of runtime, executed across multiple cores on local high-performance computing architectures. Mesh-independent results in our target metrics have been observed under the conditions above and reported in the sections to follow. All associated code, including setup scripts and post-processing functionality, are provided in the form of a GitHub repository at https://github.com/rcsc-group/BouncingDropletsSolidSurface.
@@ 4.1 Rebound metrics
Figure 5: ($a$) Coefficient of restitution $\varepsilon$ against Weber number $W\kern-1.34995pte\,$ for drops with varying $O\kern-0.45003pth\,$. Circles denote experiments and stars represent DNS results for $B\kern-0.45003pto\,=0.02$, a typical value in experiment. Solid and dashed lines are the KM model for $B\kern-0.45003pto\,=0.02$ and $B\kern-0.45003pto\,=0$, where the black dashed line shows the inertio-capillary limit ($O\kern-0.45003pth\,=0$; $B\kern-0.45003pto\,=0$). ($b$) Dimensionless contact time $\bar{t}_{c}=t_{c}/t_{\sigma}$ against $W\kern-1.34995pte\,$ for experiments, DNS, and model predictions from ($a$). The color bar maps $O\kern-0.45003pth\,$ to color on a logarithmic scale.
@@ 4.2 Spreading and retraction
Figure 6: ($a$) Dimensionless equatorial radius $\bar{r}_{m}=r_{m}/R$ against dimensionless time $\bar{t}=t/t_{\sigma}$ for experiments (markers), DNS (dotted lines), and the KM model (solid lines) for $W\kern-1.34995pte\,=2.269$ (light yellow), $W\kern-1.34995pte\,=0.253$ (green), and $W\kern-1.34995pte\,=0.023$ (dark blue), where $B\kern-0.45003pto\,=0.02$ and $O\kern-0.45003pth\,=0.03$ for all curves. ($b$) Dimensionless contact radius $\bar{r}_{c}=r_{c}/R$ against dimensionless time $\bar{t}=t/t_{\sigma}$, where line, marker type, and color are the same as ($a$). The insets show experimental results for drops of similar $W\kern-1.34995pte\,$ ($2.5-3.5$) and $B\kern-0.45003pto\,$ ($0.16-0.24$) for $O\kern-0.45003pth\,=$ 0.03, 0.08, 0.29, and 0.7, where marker darkness indicates higher $O\kern-0.45003pth\,$.
@@
The shift in rebound behavior between low and high $W\kern-1.49994pte\,$ revealed in §$\S$ 4.1 coincides with a shift in the spreading and retraction dynamics. Figure 6($a$) shows the non-dimensional equatorial radius $\bar{r}_{m}=r_{m}/R$ against non-dimensional time $\bar{t}=t/t_{\sigma}$ during contact for $W\kern-1.49994pte\,=0.023$ (dark blue), $W\kern-1.49994pte\,=0.253$ (green), and $W\kern-1.49994pte\,=2.269$ (light yellow), with $B\kern-0.50003pto\,=0.02$ and $O\kern-0.50003pth\,=0.03$. Circles represent experimental measurements, while dotted and solid lines denote DNS and the KM model results, respectively. In all cases, inertia drives spreading, causing $\bar{r}_{m}$ to increase after an initial period of localized deformation where $\bar{r}_{m}\approx 1$. As $W\kern-1.49994pte\,$ increases, the drop expands quicker and $\bar{r}_{m}$ reaches a higher maximum. After the drop reaches its maximum extent, surface tension drives retraction, decreasing $\bar{r}_{m}$. For $W\kern-1.49994pte\,=0.023$, retraction mirrors spreading, and $\bar{r}_{m}$ evolves symmetrically, consistent with quasi-static-like behavior (Moláček & Bush, 2012; Chevy et al., 2012). When $W\kern-1.49994pte\,=2.269$, the two phases are dissimilar, and $\bar{r}_{m}$ overshoots $\bar{r}_{m}=1$ during retraction, leading to apparent underdamped oscillations. The inset plot shows the evolution of $\bar{r}_{m}$ for drops of similar $W\kern-1.49994pte\,$ and $B\kern-0.50003pto\,$ but varying $O\kern-0.50003pth\,$, where darker marker shade indicates higher $O\kern-0.50003pth\,$. Increasing viscous dissipation can eliminate the appearance of underdamped behavior; however, it is not sufficient to recover the symmetry between spreading and retraction. Instead, higher $O\kern-0.50003pth\,$ shortens the spreading phase, reduces the maximum $\bar{r}_{m}$, and lengthens the retraction phase, thereby increasing asymmetry in the spreading and retraction process.
@@
The increased asymmetry between spreading and retraction as $W\kern-1.49994pte\,$ increases is also observed in the nondimensional contact radius $\bar{r}_{c}=r_{c}/R$. Figure 6($b$) shows $\bar{r}_{c}=r_{c}/R$ against nondimensional time $\bar{t}=t/t_{\sigma}$ for $W\kern-1.49994pte\,=0.023$ (dark blue), $W\kern-1.49994pte\,=0.253$ (green), and $W\kern-1.49994pte\,=2.269$ (light yellow). The data and model predictions correspond with those in figure 6($a$); thus, plot markers, line types, and colors hold the same meaning. Unlike $\bar{r}_{m}$, there is no delay in the response of $\bar{r}_{c}$ to impact as $\bar{t}=0$ is defined as the moment the drop and substrate “contact”. As $W\kern-1.49994pte\,$ increases, the maximum $\bar{r}_{c}$ increases, while the time to reach it decreases. For $W\kern-1.49994pte\,=0.023$, the spreading and retraction of $\bar{r}_{c}$ are symmetric, whereas at $W\kern-1.49994pte\,=2.269$, they are asymmetric, qualitatively resembling the behavior of $\bar{r}_{m}$ during rebound. The inset shows the evolution of $\bar{r}_{c}$ for drops of similar $B\kern-0.50003pto\,$ and $W\kern-1.49994pte\,$, but increasing $O\kern-0.50003pth\,$, as indicated by darker shaded markers. Similar to $\bar{r}_{m}$, increasing $O\kern-0.50003pth\,$ shortens the spreading phase by reducing the maximum $\bar{r}_{c}$, and extending the retraction phase. The asymmetry between spreading and retraction corresponds with the different rebound behaviors explored earlier and motivates a detailed look at their relationship across all experiments.
@@
Figure 7: The relative spreading time $t_{s}/t_{c}$ against Weber number $W\kern-1.34995pte\,$ for drops with varying $O\kern-0.45003pth\,$. Circles represent experiment and stars denote DNS for $B\kern-0.45003pto\,=0.02$. Solid lines are predictions from the KM model for $B\kern-0.45003pto\,=0.02$, and the black dashed line shows the inertio-capillary limit ($O\kern-0.45003pth\,=0$; $B\kern-0.45003pto\,=0$). The inset shows the predicted dimensionless contact radius $\bar{r}_{c}$ against dimensionless time $\bar{t}=t/t_{\sigma}$ for $W\kern-1.34995pte\,=0.04$ and $W\kern-1.34995pte\,=0.05$, where $B\kern-0.45003pto\,=0$ and $O\kern-0.45003pth\,=0$. The black dots show when $\bar{r}_{c}$ is maximum.
@@
The spreading time $t_{s}$ is the time from when the drop first contacts the substrate until $r_{c}$ reaches its maximum value $R_{c}$, and is proportional to the retraction time $t_{r}$ at low $W\kern-1.49994pte\,$, which coincides with the onset of the low-$W\kern-1.49994pte\,$ trends for $\varepsilon$ (cf. Figure 5($a$)) for each $O\kern-0.50003pth\,$. In figure 7 we show the relative spreading time $t_{s}/t_{c}$, where the contact time is the sum of the spreading and retraction time $t_{c}=t_{s}+t_{r}$, against $W\kern-1.49994pte\,$ for all $O\kern-0.50003pth\,$. Solid lines represent predictions from the KM model for $B\kern-0.50003pto\,=0.02$, with the black dashed curve representing the inertio-capillary limit ($B\kern-0.50003pto\,=0$; $O\kern-0.50003pth\,=0$). When $W\kern-1.49994pte\,\gtrapprox 0.5$, drops spread faster than they retract, and increasing $W\kern-1.49994pte\,$ increases the difference between $t_{s}$ and $t_{r}$. However, the time the drop spends spreading relative to retracting is nearly $W\kern-1.49994pte\,$-independent for $W\kern-1.49994pte\,\lessapprox 0.5$, until the drop ceases to bounce. This trend holds for all viscosities, with $O\kern-0.50003pth\,$ limiting the maximum of $t_{s}/t_{c}$. Although this maximum approaches 0.5 as $W\kern-1.49994pte\,\rightarrow 0$ in the inertio-capillary limit, it surpasses it and peaks near $W\kern-1.49994pte\,=0.04$. The inset in figure 5 shows the evolution of the dimensionless contact radius $\bar{r}_{c}$ near its maximum, which is denoted by a black dot. Oscillations due to capillary waves cause an apparent jumpiness in $t_{s}$ at low $O\kern-0.50003pth\,$. In particular, as $W\kern-1.49994pte\,$ decreases from $W\kern-1.49994pte\,=0.05$ to $W\kern-1.49994pte\,=0.04$, the maximum contact radius suddenly switches between its two largest peaks, producing the apparent sudden rise in $t_{s}$. These results reveal that the symmetry of the spreading and retraction process coincides with nearly-elastic rebound metrics that are independent of $W\kern-1.49994pte\,$. In contrast, the emergence of asymmetry reflects the onset of $W\kern-1.49994pte\,$-dependent trends for $\varepsilon$ while increasing $O\kern-0.50003pth\,$ similarly decreases $t_{s}/t_{c}$ and $\varepsilon$ across all $W\kern-1.49994pte\,$. As in the prior rebound measurements, the KM model performs best for low $W\kern-1.49994pte\,$ where as the DNS excels at higher $W\kern-1.49994pte\,$ for the spreading and retraction metrics presented throughout this section.

### v2

@@ 2.2 Experimental protocol
To begin a set of experiments, a reservoir connected to the drop generator is filled with silicone oil, and the system is manually purged to remove any trapped air (Harris et al., 2015). Next, the pulse width controlling the piezoelectric disks actuation is tuned to produce a single drop with minimal surface oscillations at impact. These settings vary based on the working fluid and nozzle. Next, the nozzle is lowered until it rests just above the substrate, and the system height is set to zero. The nozzle is raised to the desired initial drop height, typically 1 or 2 mm, and a droplet is generated. The drop falls under the action of gravity and its motion is recorded from release until bouncing ceases. After coming to rest, an external pulse of air is applied and sweeps the drop from the substrate before it can coalesce with the film, ensuring the surface remains pristine for subsequent experiments. In the few cases where the drop merged with the viscous film, the slide was translated so the next impact occurred sufficiently far from the merge site. The nozzle height is raised incrementally, and three to eight drops are recorded at each height. The experiments conclude once impact stresses become strong enough to pierce the intervening air layer (Hao et al., 2015), which occurs at $W\kern-1.49994pte\,\approx 10$ for all $O\kern-0.50003pth\,$, or drop retraction leads to droplet ejection for low $O\kern-0.50003pth\,$ drops (Richard & Quéré, 2000; Bartolo et al., 2006), as shown in supplemental movie 2.
@@ 3.1.1 Drop motion
Here, $P_{l}$ are Legendre polynomials, and $\mathscr{A^{\prime}}_{l}$ denotes the time-dependent amplitude of the $l^{\text{th}}$ mode. The $l=0$ mode is excluded due to incompressibility. The $l=1$ mode is also excluded in the present formulation due to the choice of the drop’s center of mass as the origin, since the $l=1$ mode shifts the center of mass. Figure 3($c$) demonstrates the synthesis of Legendre polynomials for modes l=24l=2-4 into a sample deformed drop shape. Similarly, the contact pressure $p^{\prime}$ exerted by the substrate on the drop is expressed as
@@
Figure 3: ($a$) Schematic of an axisymmetric drop impacting a rigid substrate, where $\zeta^{\prime}\left(\theta,t\right)$ is the dimensional interface, $h^{\prime}\left(t\right)$ the height of the center of mass of the drop, and $S\left(t\right)$ the contact surface. ($b$) Angular discretization of the right hemisphere of an axisymmetric drop during impact. Filled circles represent mesh points where the drop and substrate touch, while hollow circles denote points along the liquid-gas interface. ($c$) An example of shape reconstruction through Legendre synthesis for modes l=24l=2-4.
@@ 3.1.3 Non-dimensionalization
$\sum\limits_{l=0}^{\infty}{\mathscr{B}_{l}P_{l}(\cos(\theta))}=0,\qquad\hskip 39.83385pt\text{ for }\theta>\theta_{c};$ (27)
@@ 3.2 Direct numerical simulation
We develop a computational framework based on the Basilisk open-source code structure (Popinet, 2009, 2015) as a comparative benchmark for the mathematical model described in Section 3.1, and to interrogate quantities difficult to visualize experimentally. The fully nonlinear implementation is used to resolve the time-dependent Navier-Stokes equations in both liquid and gas phases, with physical properties of the latter taken to be the values of air at room temperature i.e. $\rho_{g}=0.00121$ g cm-3 and $\mu_{g}=0.000181$ g cm-1 s-1, in an attempt to replicate the experimental conditions. Basilisk employs the volume-of-fluid method (Hirt & Nichols (1981); Scardovelli & Zaleski (1999)) for the implicit representation of the interface describing the transition between the different fluid regions, with a so-called color function $f$ taking the value $1$ inside the liquid and 0 inside the gas regions, respectively, and values $0<f<1$ representing computational cells cut by a fluid-fluid interface. The adaptive mesh refinement and parallelization capabilities of the implementation make it a particularly good candidate for this multi-scale impact problem. The code base for bouncing problems has been constructed and systematically validated for several years in our extended group, covering fluid-structure interaction (Galeano-Rios et al., 2021) and liquid-liquid impact (Alventosa et al., 2023), with comprehensive numerical campaigns in the wider community (Sanjay et al., 2023a, b) providing key insights into related regimes.
@@
Figure 4: ($a$) Visualization of the computational domain comprising of liquid (blue) and gas regions (white), and underlying adaptive grid cell structure in the direct numerical simulation, for a typical $W\kern-1.38742pte\,=0.253$ impact taken at the point of maximal radius deformation $t=0.746$ ms. ($b$) Dimensionless distance above the substrate $\bar{z}=z/R$ against time $\bar{t}=t/t_{\sigma}$ for the top, bottom, and center of mass of a drop for $W\kern-1.38742pte\,=0.253$, $B\kern-0.46252pto\,=0.02$ and $O\kern-0.46252pth\,=0.03$. Black solid lines represents experiment, while blue solid lines and yellow dashed lines represent the KM model and direct numerical simulations, respectively. ($c$) Overlay of the predicted shape from the KM model (solid blue) and direct numerical simulations (dashed yellow) with experimental images corresponding to ($b$). The drop radius is $R=0.203$ mm and the time interval between images is 0.375 ms.
[…]
@@ 4.1 Rebound metrics
Figure 5: ($a$) Coefficient of restitution $\varepsilon$ against Weber number $W\kern-1.38742pte\,$ for drops with varying $O\kern-0.46252pth\,$. Circles denote experiments and stars represent DNS results for $B\kern-0.46252pto\,=0.02$, a typical value in experiment. Solid and dashed lines are the KM model for $B\kern-0.46252pto\,=0.02$ and $B\kern-0.46252pto\,=0$, where the black dashed line shows the inertio-capillary limit ($O\kern-0.46252pth\,=0$; $B\kern-0.46252pto\,=0$). ($b$) Dimensionless contact time $\bar{t}_{c}=t_{c}/t_{\sigma}$ against $W\kern-1.38742pte\,$ for experiments, DNS, and model predictions from ($a$). The color bar maps $O\kern-0.46252pth\,$ to color on a logarithmic scale.
@@ 4.2 Spreading and retraction
Figure 6: ($a$) Dimensionless equatorial radius $\bar{r}_{m}=r_{m}/R$ against dimensionless time $\bar{t}=t/t_{\sigma}$ for experiments (markers), DNS (dotted lines), and the KM model (solid lines) for $W\kern-1.38742pte\,=2.269$ (light yellow), $W\kern-1.38742pte\,=0.253$ (green), and $W\kern-1.38742pte\,=0.023$ (dark blue), where $B\kern-0.46252pto\,=0.02$ and $O\kern-0.46252pth\,=0.03$ for all curves. ($b$) Dimensionless contact radius $\bar{r}_{c}=r_{c}/R$ against dimensionless time $\bar{t}=t/t_{\sigma}$, where line, marker type, and color are the same as ($a$). The insets show experimental results for drops of similar $W\kern-1.38742pte\,$ (2.53.52.5-3.5) and $B\kern-0.46252pto\,$ (0.160.240.16-0.24) for $O\kern-0.46252pth\,=$ 0.03, 0.08, 0.29, and 0.7, where marker darkness indicates higher $O\kern-0.46252pth\,$.
@@
The shift in rebound behavior between low and high $W\kern-1.49994pte\,$ revealed in $\S$ 4.1 coincides with a shift in the spreading and retraction dynamics. Figure 6($a$) shows the non-dimensional equatorial radius $\bar{r}_{m}=r_{m}/R$ against non-dimensional time $\bar{t}=t/t_{\sigma}$ during contact for $W\kern-1.49994pte\,=0.023$ (dark blue), $W\kern-1.49994pte\,=0.253$ (green), and $W\kern-1.49994pte\,=2.269$ (light yellow), with $B\kern-0.50003pto\,=0.02$ and $O\kern-0.50003pth\,=0.03$. Circles represent experimental measurements, while dotted and solid lines denote DNS and the KM model results, respectively. In all cases, inertia drives spreading, causing $\bar{r}_{m}$ to increase after an initial period of localized deformation where $\bar{r}_{m}\approx 1$. As $W\kern-1.49994pte\,$ increases, the drop expands quicker and $\bar{r}_{m}$ reaches a higher maximum. After the drop reaches its maximum extent, surface tension drives retraction, decreasing $\bar{r}_{m}$. For $W\kern-1.49994pte\,=0.023$, retraction mirrors spreading, and $\bar{r}_{m}$ evolves symmetrically, consistent with quasi-static-like behavior (Moláček & Bush, 2012; Chevy et al., 2012). When $W\kern-1.49994pte\,=2.269$, the two phases are dissimilar, and $\bar{r}_{m}$ overshoots $\bar{r}_{m}=1$ during retraction, leading to apparent underdamped oscillations. The inset plot shows the evolution of $\bar{r}_{m}$ for drops of similar $W\kern-1.49994pte\,$ and $B\kern-0.50003pto\,$ but varying $O\kern-0.50003pth\,$, where darker marker shade indicates higher $O\kern-0.50003pth\,$. Increasing viscous dissipation can eliminate the appearance of underdamped behavior; however, it is not sufficient to recover the symmetry between spreading and retraction. Instead, higher $O\kern-0.50003pth\,$ shortens the spreading phase, reduces the maximum $\bar{r}_{m}$, and lengthens the retraction phase, thereby increasing asymmetry in the spreading and retraction process.
@@
The increased asymmetry between spreading and retraction as $W\kern-1.49994pte\,$ increases is also observed in the nondimensional contact radius $\bar{r}_{c}=r_{c}/R$. Figure 6($b$) shows $\bar{r}_{c}=r_{c}/R$ against nondimensional time $\bar{t}=t/t_{\sigma}$ for $W\kern-1.49994pte\,=0.023$ (dark blue), $W\kern-1.49994pte\,=0.253$ (green), and $W\kern-1.49994pte\,=2.269$ (light yellow). The data and model predictions correspond with those in figure 6($a$); thus, plot markers, line types, and colors hold the same meaning. Unlike $\bar{r}_{m}$, there is no delay in the response of $\bar{r}_{c}$ to impact as $\bar{t}=0$ is defined as the moment the drop and substrate “contact”. As $W\kern-1.49994pte\,$ increases, the maximum $\bar{r}_{c}$ increases, while the time to reach it decreases. For $W\kern-1.49994pte\,=0.023$, the spreading and retraction of $\bar{r}_{c}$ are symmetric, whereas at $W\kern-1.49994pte\,=2.269$, they are asymmetric, qualitatively resembling the behavior of $\bar{r}_{m}$ during rebound. The inset shows the evolution of $\bar{r}_{c}$ for drops of similar $B\kern-0.50003pto\,$ and $W\kern-1.49994pte\,$, but increasing $O\kern-0.50003pth\,$, as indicated by darker shaded markers. Similar to $\bar{r}_{m}$, increasing $O\kern-0.50003pth\,$ shortens the spreading phase by reducing the maximum $\bar{r}_{c}$, and extending the retraction phase. The asymmetry between spreading and retraction motivates a detailed look at their relationship across all experiments.
@@
Figure 7: The relative spreading time $t_{s}/t_{c}$ against Weber number $W\kern-1.38742pte\,$ for drops with varying $O\kern-0.46252pth\,$. Circles represent experiment and stars denote DNS for $B\kern-0.46252pto\,=0.02$. Solid lines are predictions from the KM model for $B\kern-0.46252pto\,=0.02$, and the black dashed line shows the inertio-capillary limit ($O\kern-0.46252pth\,=0$; $B\kern-0.46252pto\,=0$). The inset shows the predicted dimensionless contact radius $\bar{r}_{c}$ against dimensionless time $\bar{t}=t/t_{\sigma}$ for $W\kern-1.38742pte\,=0.04$ and $W\kern-1.38742pte\,=0.05$, where $B\kern-0.46252pto\,=0$ and $O\kern-0.46252pth\,=0$. The black dots show when $\bar{r}_{c}$ is maximum.
@@
The spreading time $t_{s}$ is the time from when the drop first contacts the substrate until $r_{c}$ reaches its maximum value $R_{c}$, and is proportional to the retraction time $t_{r}$ at low $W\kern-1.49994pte\,$. In figure 7 we show the relative spreading time $t_{s}/t_{c}$, where the contact time is the sum of the spreading and retraction time $t_{c}=t_{s}+t_{r}$, against $W\kern-1.49994pte\,$ for all $O\kern-0.50003pth\,$. Solid lines represent predictions from the KM model for $B\kern-0.50003pto\,=0.02$, with the black dashed curve representing the inertio-capillary limit ($B\kern-0.50003pto\,=0$; $O\kern-0.50003pth\,=0$). When $W\kern-1.49994pte\,\gtrapprox 0.5$, drops spread faster than they retract, and increasing $W\kern-1.49994pte\,$ increases the difference between $t_{s}$ and $t_{r}$. However, the time the drop spends spreading relative to retracting is nearly $W\kern-1.49994pte\,$-independent for $W\kern-1.49994pte\,\lessapprox 0.5$, until the drop ceases to bounce. This trend holds for all viscosities, with $O\kern-0.50003pth\,$ limiting the maximum of $t_{s}/t_{c}$. Although this maximum approaches 0.5 as $W\kern-1.49994pte\,\rightarrow 0$ in the inertio-capillary limit, it surpasses it and peaks near $W\kern-1.49994pte\,=0.04$. The inset in figure 5 shows the evolution of the dimensionless contact radius $\bar{r}_{c}$ near its maximum, which is denoted by a black dot. Oscillations due to capillary waves cause an apparent jumpiness in $t_{s}$ at low $O\kern-0.50003pth\,$. In particular, as $W\kern-1.49994pte\,$ decreases from $W\kern-1.49994pte\,=0.05$ to $W\kern-1.49994pte\,=0.04$, the maximum contact radius suddenly switches between its two largest peaks, producing the apparent sudden rise in $t_{s}$. These results reveal that the symmetry of the spreading and retraction process coincides with nearly-elastic rebound metrics that are independent of $W\kern-1.49994pte\,$ in the inertio-capillary limit. In contrast, the emergence of asymmetry coincides with the onset of $W\kern-1.49994pte\,$-dependent trends for $\varepsilon$ while increasing $O\kern-0.50003pth\,$ similarly decreases $t_{s}/t_{c}$ and $\varepsilon$ across all $W\kern-1.49994pte\,$. As in the prior rebound measurements, the KM model performs best for low $W\kern-1.49994pte\,$ where as the DNS excels at higher $W\kern-1.49994pte\,$ for the spreading and retraction metrics presented throughout this section.